# Comparison of Generative Learning Methods for Turbulence Modeling

### v1

@@ 6 Conclusion and Outlook
We compared three generative learning approaches, starting from a VAE, a lighter model with about 4 million parameters, to DCGAN, already widely studied in the field of neural network turbulence modeling, to DDPM, a recent generative model. Our experiments have shown that the VAE is not able to match the quality of the LES, which can be directly observed by just taking a look at the generated image. In contrast, DCGAN and DDPM both synthesize turbulent flow that agrees excellently with the LES flow at the visual level and by examining physics-based metrics. Nevertheless, we found that DCGAN performs better in hitting the peaks of the distributions regarding the physics-based evaluation. Finally, a strong argument for DCGAN was found by analyzing the computational cost, since it requires significantly less data to train and is significantly faster in training and inference time compared to DDPM. In conclusion, figure 9 shows that DCGAN is clearly superior to DDPM, especially in terms of model efficiency. Our experiments have shown that the DCGAN itself, as well as extended variations based on the DCGAN, remain promising for future work in turbulence modeling. However, with the growing amount of data and increasingly powerful computers, DDPM should not be completely discarded for this problem. They are not as efficient as DCGAN, but they are still significantly faster in generating turbulent flow than LES, and the quality is also adequate.
@@ Acknowledgments
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) within the priority program "Carnot Batteries: Inverse Design from Markets to Molecules" (SPP 2403) grant no. 526152410.

### v2

@@ 6 Conclusion and Outlook
We compared three generative learning approaches, starting from a VAE, a lighter model with about 4 million parameters, to DCGAN, already widely studied in the field of neural network turbulence modeling, to DDPM, a recent generative model. Our experiments have shown that the VAE is not able to match the quality of the LES, which can be directly observed by just taking a look at the generated image. In contrast, DCGAN and DDPM both synthesize turbulent flow that agrees excellently with the LES flow at the visual level and by examining physics-based metrics. Nevertheless, we found that DCGAN performs better in hitting the peaks of the distributions regarding the physics-based evaluation, but care has to be taken to reduce noise in experimental data. This issue is particularly relevant for the experimental PIV data. While computational cost plays a minor role, both DCGAN and DDPM accurately reproduce the measured flow statistics and small-scale structures, despite the inherent measurement uncertainties. In particular, the DCGAN captures the characteristic small-scale fluctuations of the PIV data, demonstrating that these models can robustly handle experimental noise and variability. Returning to the overall comparison, a strong argument for DCGAN was found by analyzing the computational cost, since it requires significantly less data to train and is significantly faster in training and inference time compared to DDPM. In conclusion, both figure˜14 and figure˜15 show that DCGAN is clearly superior to DDPM, especially in terms of model efficiency. Our experiments have shown that the DCGAN itself, as well as extended variations based on the DCGAN, remain promising for future work in turbulence modeling. However, with the growing amount of data and increasingly powerful computers, DDPM should not be completely discarded for this problem. They are not as efficient as DCGAN, but they are still significantly faster in generating turbulent flow than LES, and the quality is also adequate. Overall, our experiments indicate that DCGAN and its extended variations remain highly promising for future work in turbulence modeling, including both numerical simulations and experimental datasets. The successful application to PIV data suggests that these AI-based approaches are capable not only of reproducing LES-like flows but also of accommodating the uncertainties inherent in experimental measurements, broadening their potential impact in fluid dynamics research.
@@ Acknowledgments
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) within the priority program "Carnot Batteries: Inverse Design from Markets to Molecules" (SPP 2403) grant no. 526152410 and 525893212 and DFG grant no. 467227170. We also thank the two anonymous referees for hints that helped to improve this paper.
Appendix A Evaluation results for larger grid of pixels
A.1 Flow around a cylinder
13D
16D
20D
Figure 16: Comparison of the mean local velocity fluctuation magnitude was performed over increasingly large portions of the downstream wake region, extending to $13D$, $16D$, and $20D$ within the numerical domain behind the cylinder (see figure˜6(a)), corresponding to $50\%$, $75\%$, and $100\%$ of the investigated wake. Each value represents the deviation of the local flow velocity from the background (mean) flow. All datasets were normalized prior to evaluation.
13D
16D
20D
Figure 17: Comparison of the variance of the local velocity fluctuation was performed over increasingly large portions of the downstream wake region, extending to $13D$, $16D$, and $20D$ within the numerical domain behind the cylinder (see figure˜6(a)), corresponding to $50\%$, $75\%$, and $100\%$ of the investigated wake. Each value represents the deviation of the local flow velocity from the background (mean) flow. All datasets were normalized prior to evaluation.
A.2 Flow behind array of seven cylinders
25.2D
37.5D
50D
Figure 18: Comparison of the mean local velocity fluctuation magnitude was performed over increasingly large portions of the downstream wake region, extending to $25.2D$, $37.5D$, and $50D$ within the numerical domain behind the array of seven cylinders in the 4V2H test case (see figure˜6(b)), corresponding to $50\%$, $75\%$, and $100\%$ of the investigated wake. Each value represents the deviation of the local flow velocity from the background (mean) flow. All datasets were normalized prior to evaluation.
25.2D
37.5D
50D
Figure 19: Comparison of the variance of the local velocity fluctuation was performed over increasingly large portions of the downstream wake region, extending to $25.2D$, $37.5D$, and $50D$ within the numerical domain behind the array of seven cylinders in the 4V2H test case (see figure˜6(b)), corresponding to $50\%$, $75\%$, and $100\%$ of the investigated wake. Each value represents the deviation of the local flow velocity from the background (mean) flow. All datasets were normalized prior to evaluation.
25.2D
37.5D
50D
Figure 20: Comparison of the mean local velocity fluctuation magnitude was performed over increasingly large portions of the downstream wake region, extending to $25.2D$, $37.5D$, and $50D$ within the numerical domain behind the array of seven cylinders in the 2V8H test case (see figure˜6(b)), corresponding to $50\%$, $75\%$, and $100\%$ of the investigated wake. Each value represents the deviation of the local flow velocity from the background (mean) flow. All datasets were normalized prior to evaluation.
25.2D
37.5D
50D
Figure 21: Comparison of the variance of the local velocity fluctuation was performed over increasingly large portions of the downstream wake region, extending to $25.2D$, $37.5D$, and $50D$ within the numerical domain behind the array of seven cylinders in the 2V8H test case (see figure˜6(b)), corresponding to $50\%$, $75\%$, and $100\%$ of the investigated wake. Each value represents the deviation of the local flow velocity from the background (mean) flow. All datasets were normalized prior to evaluation.